OpenStudy (sissyedgar):

Math for Meh! :P What is the solution to the equation? 6x – 27 = –3 A. x = –5 B. x = –4 C. x = 4 D. x = 5

OpenStudy (anonymous):

Add 27 to both sides to cancel it from the left side so it'll look like 6x=24, then solve for x.

OpenStudy (sissyedgar):

hat is the solution to the equation? 2q + 18 = –5q – 3 A. q = –5 B. q = –3 C. q = 3 D. q = 5

OpenStudy (misty1212):

add \(-5q\) get \[7q+18=-3\] subtract \(18\) get \[7q=-21\] divide by \(7\)

OpenStudy (sissyedgar):

B.

OpenStudy (misty1212):

yup

OpenStudy (sissyedgar):

The product of 3 and a number is 12 more than the sum of the number and 26. What is the number?

OpenStudy (misty1212):

\[3x=x+26+12\]

OpenStudy (sissyedgar):

mhm...

OpenStudy (misty1212):

so \[2x=38\]

OpenStudy (sissyedgar):

divided by 2?

OpenStudy (misty1212):

yes

OpenStudy (sissyedgar):

19

OpenStudy (misty1212):

yes